A concert sells out a 20,000 seat arena. Main seat tickets cost $55, and back seat tickets for $45. The concert made $955,000. How many main seat tickets and back seat tickets were sold?
Answer by TutorDelphia(193)

We know the total seats which are made up of main and back seats equals 20,000:
m+b=20,000
We know the earnings were 55 for every main and 45 for every back:
55*m+45*b=955,000
We now have a system of equations to solve. Lets multiply both sides of the first equation by 45

Now subtract :
45m+45b=900,000 from
55*m+45*b=955,000
That gives us
10m=55,000
so
m=5,500
And if we sub m back into our first equation we find that
b=14,500